simplify (-2-2)[8+(7+8)]
Answer by lenny460(1073)

(-2 - 2)[8 + (7 + 8)]
(-2 - 2) = -2 - 2 = -4
-4[8 + (7 + 8)]
(7 + 8) = 7 + 8 = 15
-4[8 + 15]
[8 + 15] = 8 + 15 = 23
-4[23] = -4 * 23 = -92
The Answer:
-92